# AgilePilot: DRL-Based Drone Agent for Real-Time Motion Planning in Dynamic Environments by Leveraging Object Detection

### v1

@@ V-A Experimental Setup
Fig. 10 illustrates the general layout of our experimental setup. The drone, with a mass of 1.80 kg, is equipped with a high-performance Intel NUC onboard computer, alongside a RealSense T265 camera for real-time object detection. The onboard computer communicates with a SpeedyBee flight controller, which is responsible for sending and receiving commands through the ArduPilot firmware. For environmental layout, obstacle, gate and landing pad are used where the objects are dynamically moved manually by a rope at speeds between 0.4-0.7 m/s to validate the transfer of simulation results to real-world conditions.
@@
Case 1: Slow left moving gate with obstacle and target.
Case 2: Fast right moving gate with obstacle and target.
@@ V-B2 Action predictions by DRL agent
The position and velocity plots in Fig. 12 and Fig. 13 summarises that in Case 1, the gate was moved at a steady pace, allowing the drone to navigate without any abrupt changes in its motion. The agent, recognising that the environment contained moving objects, gradually increased its speed to a maximum of 1.5 m/s while avoding the obstacle smoothly. This smooth velocity profile suggests that the agent anticipated the gate’s motion effectively, managing to navigate without encountering any significant difficulties.
In Case 2, the gate moved much faster as the drone approached it, presenting a more dynamic scenario. To adjust, the drone increased its velocity up to 3.0 m/s in response to the rapid movement of the gate. This quick adjustment also demonstrated the gate detection system’s ability to accurately track the fast-moving gate. This behaviour specifically underscores the robustness of the agent’s velocity management in dynamic conditions.
@@ VI Conclusion and Future work
In conclusion, we presented AgilePilot, a novel DRL-based motion planner that enables intelligent navigation in dynamic environments by leveraging real-time object detection. The DRL framework was trained in dynamic simulation environments, facilitating a robust and adaptive learning process that bridges the simulation-to-reality gap effectively. The performance of our motion planner was compared with an APF-based motion planner in simulated dynamic environments. In summary, AgilePilot outperforms the classical approach by completing the task times faster and also times with more accuracy than APF planner in terms of dynamic target tracking. Overall, our trained agent exhibited an average success rate of out of a total 75 conducted experiments.
In the real world, three conditions were tested with varying environment and dynamic conditions by deploying the trained DRL agent onto hardware that utilized real-time object detection. The drone demonstrated safe navigation while avoiding obstacles and steadily passing through dynamic gate. However, when the environment was subject to high-speed changes, such as rapidly moving gates, the drone displayed remarkable agility and adaptability, adjusting its speed up to 3 m/s to meet the dynamic demands of the environment. The real-time detection of gates and obstacles was accurate by encompassing RMSE of 0.076 m for obstacles and 0.37 m for the moving gates. The observed delay results from the time taken for image acquisition, neural network processing, and the transformation of position and pose filtering. However, this delay does not significantly impact performance, and the drone performs well in experiments. These results validate the effectiveness of AgilePilot in real-world conditions, demonstrating its capability to navigate dynamic and unpredictable environments using real-time velocity predictions from trained agent.

### v2

@@ V-A Experimental Setup
Fig. 10 shows the general layout of our experimental setup. The drone, with a mass of 1.80 kg, is equipped with a high-performance Intel NUC onboard computer, alongside a RealSense T265 camera for real-time object detection. The onboard computer communicates with a SpeedyBee flight controller, which is responsible for sending and receiving commands through the ArduPilot firmware. For environmental layout, obstacles, gates, and landing pads are used where the objects are dynamically moved manually by a rope at speeds between 0.4 and 0.7 m/s to validate the transfer of simulation results to real-world conditions.
@@
Case 1: Slow left-moving gate with obstacle and target.
Case 2: Fast right-moving gate with obstacle and target.
@@ V-B2 Action predictions by DRL agent
The position and velocity plots in Fig. 12 and Fig. 13 summarize that in Case 1 the drone was able to navigate without any abrupt changes through the gate being moved at a constant speed. The agent detects moving objects and gradually increases its speed to a maximum of 1.5 m/s while avoiding the obstacle smoothly. This smooth velocity profile suggests that the agent anticipated the gate’s motion effectively.
In Case 2, the gate moved much faster as the drone approached it, presenting a highly dynamic scenario. To adjust, the drone increased its velocity up to 3.0 m/s. This quick adjustment also demonstrated the gate detection system’s ability to accurately track the fast-moving gate. This behavior specifically underscores the robustness of the agent’s velocity management in dynamic conditions.
@@ VI Conclusion and Future work
In conclusion, we presented AgilePilot, a novel DRL-based motion planner that enables intelligent navigation in dynamic environments by leveraging real-time object detection. The DRL framework was trained in dynamic simulation environments, facilitating a robust and adaptive learning process that bridges the simulation-to-reality gap effectively. The performance of our motion planner was compared with an APF-based motion planner in simulated dynamic environments. In summary, AgilePilot outperforms the classical approach by completing the task times faster and also times with more accuracy than the APF planner in terms of dynamic target tracking. Overall, our trained agent exhibited an average success rate of out of a total of 75 conducted experiments.
In the real world, three conditions were tested with varying environments and dynamic conditions by deploying the trained DRL agent onto hardware that utilized real-time object detection. The drone demonstrated safe navigation while avoiding obstacles and steadily passing through the dynamic gate. However, when the environment was subject to high-speed changes, such as rapidly moving gates, the drone displayed remarkable agility and adaptability, adjusting its speed up to 3 m/s to meet the dynamic demands of the environment. The real-time detection of gates and obstacles was accurate, encompassing an RMSE of 0.076 m for obstacles and 0.37 m for the moving gates. The observed delay results from the time taken for image acquisition, neural network processing, and the transformation of position and pose filtering. However, this delay does not significantly impact performance, and the drone performs well in experiments. These results validate the effectiveness of AgilePilot in real-world conditions, demonstrating its capability to navigate dynamic and unpredictable environments using real-time velocity predictions from trained agents.
@@
Acknowledgements
Research reported in this publication was financially supported by the RSF grant No. 24-41-02039.